Circular No. 6101

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

GS 1826–24 C. Motch, Observatoire de Strasbourg; D. Barret, Centre d’Etude

Spatiale des Rayonnements and Harvard-Smithsonian Center for Astrophysics; W. Pietsch, Max-Planck-Institut für Extraterrestrische Physik; G. Hasinger, Astrophysikalisches Institut, Potsdam; and E. Giraud, European Southern Observatory (ESO), write: “We report the optical identification of the black-hole candidate and x-ray source GS 1826–24 (Tanaka 1989, ESA SP-296, p. 3) with a faint (V = 19, B−V = +0.4, U −B = −0.5) ultravioletexcess object exhibiting the He II 468.6-nm (equivalent width = −0.4 nm) and Hα (EW = −0.18 nm) lines in emission. Optical observations were obtained with the ESO New Technology Telescope (+ EMMI) on June 3–4 and Oct. 20. The proposed optical counterpart is located at α = 18^{h}29^{m}28^{s}.2, δ = −23^{o}47′49′′.1 (equinox 2000.0; ±2′′), well within the ROSAT PSPC pointing error circle (90-percent confidence radius 18′′) centered at position end figures 27^{s}.63, 53′′.5 (see also Johnston et al. 1994, A.Ap. 289, 763). The x-ray source was detected during the ROSAT All-Sky Survey and subsequent 20000-s pointing on 1992 Oct. 9–12, with a mean PSPC countrate of 4 and 2 counts/s, respectively. Our optical identification indicates that the black-hole candidate is accreting from a low-mass companion. Finding charts may be provided upon request to motch@cdsxb7.u-strasbg.fr.”

X-RAY NOVA IN SCORPIUS S. N. Zhang, B. A. Harmon, W. S. Paciesas, C. A. Wilson, and G. J.

Fishman report for the Compton Gamma Ray Observatory BATSE Team: “The x-ray nova in Scorpius (first reported on IAUC 6046) is again in outburst. The source intensity, in mCrab for the band 20–100 keV, was < 100 mCrab on Oct. 31, then reached 290 ± 70 on Nov. 1, 360 ± 25 on Nov. 2, 350 ± 30 on Nov. 3, 680 ± 50 on Nov. 5, 740 ± 30 on Nov. 6, and 710 ± 35 on Nov. 7. The flux during the last two days is near the peak level of the previous outburst (IAUC 6075). The photon spectrum is well fit with a power law with an index of −3.1 ± 0.2 from 20 to 100 keV. We urge immediate follow-up observations of this source at other wavelengths. It should be noted that the previous outburst only lasted for ∼10 days within the energy range 20–200 keV.”

1994 November 9 Daniel W. E. Green